OpenStudy (meilendurcer):

Please help i don't understand this.!! Complete the table. Below a) 5 6 7 8 9 b) 3.8 ,4.6, 5.4

OpenStudy (hba):

Where are we ? Any topic ? Sub topics ?

OpenStudy (hba):

Where is the table ?

OpenStudy (meilendurcer):

no, the table is at the top where A and B are

OpenStudy (hba):

So we have to complete the b part ?

OpenStudy (meilendurcer):

yes

OpenStudy (hba):

If that is the case You should check the diffrence d=4.6-3.8=0.8 Now just add 0.8 in 5.4 and then you can add 0.8 again in it. So you make a nice table.

OpenStudy (hba):

@meilendurcer You with me ? Got it ?

OpenStudy (meilendurcer):

yes

OpenStudy (hba):

Well then i guess i am done here. So bye.

OpenStudy (meilendurcer):

so 6.2? and so on?

OpenStudy (hba):

Yeah 6.2+0.8=

OpenStudy (hba):

@meilendurcer Are you done with this question ?

OpenStudy (meilendurcer):

got it and this one? Find the next three terms of the sequence. –4, –8, –16, –32, . . .

OpenStudy (meilendurcer):

A. –96, –288, –864 B. –64, –128, –256 C. –64, –288, –256 D. –36, –40, –44

OpenStudy (hba):

What is the diffrence here ?

OpenStudy (hba):

The Second term - first term \[T_2-T_1=\]

OpenStudy (hba):

@meilendurcer Are you with me ?

OpenStudy (meilendurcer):

Yea

OpenStudy (hba):

So tell me what will be T_2-T_1= Here T_2=-8 T_1=-4 @meilendurcer Please find the diffrence for me ,Please.

OpenStudy (meilendurcer):

one sec

OpenStudy (meilendurcer):

-4?

OpenStudy (hba):

Sorry to say but this is not arithmetic sequence it is geometric so you have to find the ratio Ratio=T_2/T_1 Do this for me ?

OpenStudy (hba):

@meilendurcer ?

OpenStudy (hba):

@meilendurcer I know you have a busy life but we need to solve this question,so please respond.

OpenStudy (meilendurcer):

sorry i am back

OpenStudy (meilendurcer):

@Hero

OpenStudy (meilendurcer):

Yes, i can't i get confused with numbers

OpenStudy (meilendurcer):

@BrittanyFrench

OpenStudy (anonymous):

Yes? >:)

OpenStudy (meilendurcer):

-8

OpenStudy (meilendurcer):

-16